# Towards agent-based-model informed neural networks

### v1

@@ Results
where $h_{\ell}\in\mathbb{R}^{d_{\ell}}$ denotes the activations at layer $\ell$, $e_{i}\in\mathbb{R}^{d_{e}}$ is the static embedding for country $i$, and $\odot$ is element-wise multiplication. Intuitively, the adapters rescale and shift each hidden unit so that heterogeneous countries can share the same base network while having country-specific gain and bias patterns; this greatly reduces the parameter count relative to training a separate network per economy, yet retains the ability to express qualitatively different responses to shocks. The latent macro network $\phi_{3}$ receives $y_{i},x_{i},e_{i}]$ and outputs $\dot{y}_{i}$, letting unemployment, inflation, and other channels react causally to GDP and other macro channels deviations within the same differentiable rollout. For more details on the mathematical formulation for the forward pass of the macro-economic model, detailing the element-wise dynamics for country $i$ and the FiLM adaptation mechanism, see Appendix Auxiliary Dynamics ($\phi_{y}$):.
@@
The interaction term $\psi_{\text{GDP}}$ reduces to the classical GLV form $x_{i}(\max(x_{j},0)+\varepsilon)^{\beta}$ when adapters are neutral. We learn $\beta$ jointly with $A$ (converging to 0.20.20.40.4 in practice) to attenuate spillovers from very large economies, and optionally gate each row with $\tanh(\omega_{i})$ when interaction scaling is enabled. The matrix $A$ starts from a weakly coupled baseline ($A_{ij}=1/N$, $A_{ii}=0$) and is clamped to a user-specified maximum norm to prevent runaway feedback loops.
@@
Training on 1995–2020 data followed by 2021–2024 evaluation shows that the shared-parameter GLV reproduces the held-out trajectories while learning interpretable couplings, as illustrated in Fig. 3. Because the architecture is differentiable with respect to $A$, we can implement counterfactual edits such as zeroing the bilateral US–China channels $A_{\text{US},\text{CN}}$ and $A_{\text{CN},\text{US}}$. Appendix Fig. 10 compares historical data, baseline forecasts, and the edited intervention rollout initialized from the true 2024 state, highlighting how severing a single pair of links modifies the downstream trajectory while leaving the rest of the network largely untouched.
@@ Conclusion
Empirically, we showed that a model trained on a single trajectory generalizes out-of-sample to unseen graphs, accurately reproducing aggregate epidemic curves. We further demonstrated intervention analysis: despite never observing interventions during training, the learned model reliably predicts the system’s response, with predicted curves closely matching the ground truth. In particular, we showed that the model accurately captures the response to a time-bounded social-distancing window, highlighting its utility for counterfactual scenarios and control [Bottcher2022AIPontryagin]. The key reason these models succeed in counterfactual analysis is that they learn the underlying dynamics of the system rather than mere correlations, enabling them to predict responses to novel interventions. However, the model cannot respond to arbitrary interventions; it is limited to the interventions that can be incorporated into the learned differential equations, via additive or multiplicative terms in the differential equations. We have compared the model to other graph neural networks in neural ODE setting, and showed that abm-informed neural networks can better capture the dynamics out of sample than the state-of-the-art graph neural networks. Futhermore, abm-nn model is robust to small perturbations in the data, such as observational noise, and even, it can act as a regularization term for the stability of the model. Finally, we have applied the model to the macroeconomic case study, and showed that it can learn interpretable couplings between macro variables and GDP, and can be used to simulate counterfactual scenarios by editing the interaction structure.
@@
By moving away from the theoretical purity of perfect rationality, our framework better reflects the reality of out-of-equilibrium systems through specific inductive biases. We enforce informational boundedness via restricted Graph Neural Networks, ensuring agents react only to local peers rather than global aggregates. Furthermore, the Neural ODE framework allows us to model the speed of rationality: the discretization step of the underlying solver acts as a proxy for cognitive latency, where the step size dictates how quickly agents can process and react to messages from their neighbors. This temporal constraint is matched by a computational one; the use of gradient descent calibrated against empirical data implies that agents learn ’good enough’ local optima—heuristics—rather than global solutions, mirroring human satisficing behavior. When combined with FiLM-based heterogeneity and restricted graph topologies, we replace the frictionless ’representative agent’ with a population of diverse, friction-constrained actors. Ultimately, by integrating these temporal and computational limits with conservation laws, the model offers a plausible mechanism for simulating how constrained agents navigate structural changes.
@@
To conclude, we believe that the proposed framework is not a final solution, but rather a starting point for principled, interpretable neural modeling of complex systems. Future work will explore stochastic differential equation (SDE) formulations to explicitly model process noise and uncertainty quantification, extended Kalman filtering appliactions, as well as scaling to larger heterogeneous agent populations for more realistic economic modelling.
@@
The models presented in this paper fall within the category of agent-based models coupled with neural networks. If these models are used to inform policy recommendations, they should be evaluated with extra due diligence, where all assumptions are explicitly reported together with their limitations. Models that are evaluated only in simulation environments have inherent limitations that arise from the fidelity and scope of the simulation environment itself. Real-world deployment requires careful validation against empirical data and consideration of factors not captured in the simulation.
Any model of socio-techno-economic systems, should not be used in unethical ways to harm people, countries, or organizations. Practitioners deploying these models must consider potential misuse, including discriminatory applications, manipulation of social systems, or deployment without informed consent of affected populations. Models of social systems should always be designed with the principle that, in the long run, they increase the well-being of humans. Decision-makers must ensure that model-driven interventions respect democratic principles, protect human rights, and account for distributional equity and justice.
@@ Acknowledgements
The experiment in Fig. 4 uses the command listed above with hyperparameters: $\text{epochs}=500$, base learning rate $5\times 10^{-4}$ (AdamW with cyclical max learning rate $2\times 10^{-2}$, cyclical step size =75=75), RK4 rollout horizon $t_{\max}=250$ with training truncated at $t_{\text{end}}=50$, curriculum initialized at $T_{\text{base}}=5$ and incremented by $4$ every $50$ epochs, gradient clipping at $2.0$, and stagnation threshold $50$ epochs before forcing a curriculum increase.
@@
We assess how the macro-level learner behaves when the aggregate SIR curves are corrupted by observational noise. Using the noise-robustness variant of the learning algorithm, we replay the SIR-on-graph experiment with the same Erdős–Rényi prior ($n=100$, $p=0.05$), base parameters ($\beta=0.4$, $\gamma=0.2$), curriculum schedule (1.0 horizon increment every 50 epochs starting at $t=1$), and the triangular learning-rate policy ($\text{base }lr=10^{-4}$, $\text{max }lr=5\times 10^{-4}$, step 100). We vary only the observational noise injected into $s(t),i(t),r(t)$ for $t\geq 3$: Gaussian noise with $\sigma\in\{1.0,2.0,3.0\}$ is added per time step, while the differentiable learner sees only these noisy trajectories. Each run trains for 1000–1500 epochs (longer for higher noise) and logs the full command/config in the corresponding experiment folder.
Figure 6: Noise-robustness study for the macro SIR learner. Columns correspond to observational noise $\sigma=1.0$ (left), $2.0$ (middle), and $3.0$ (right). Each panel overlays the true aggregate curves (solid), noisy observations (dotted), and learned trajectories (dashed with diamonds). Despite the increasing corruption after $t=3$, the ABM-NN continues to track the underlying dynamics and maintains separation between in-sample (denoted by the shaded training region) and out-of-sample forecasts. See the experiment directories for full configurations and commands.
@@ 3. Detailed Component Definitions
$\psi_{i}=\mathcal{C}(k)\cdot s_{i}\cdot[x_{i}]_{+}\sum_{j}A_{ij}([x_{j}]_{+}+\epsilon)^{\beta},$ (44)
@@
  • $s_{i}$ is an optional per-country scaling factor (defaults to 1.0 unless learnable interaction scales are active, in which case $s_{i}=\tanh(\omega_{i})S_{\max}$).
@@
The optimizer learns the interaction matrix $\mathbf{A}$, the exponent $\beta$, and the optional scaling logits $\omega_{i}$ (if enabled).
@@
    • FiLM adapters applied to all hidden layers as described in Eqs. (676)-(680).
@@
    • Training period: 1995–2021 (cutoff year).
@@
The macro panel used throughout the paper originates from a data utility pipeline, which queries the World Bank API for the ten largest economies (GDP, unemployment, lending interest rate, sovereign debt, working-age population share, CPI inflation, current account balance) and augments the table with the global oil price index from FRED (POILAPSPINDEXQ). The raw annual series covering 1995–2024 are cached in the data cache directory. We then apply a strictly causal constant-velocity Kalman filter to each country/indicator pair using variable-specific process/measurement noise levels; missing values are imputed forward in time without peeking at future observations and the result is stored in the processed dataset. All statistics and plots in the main text reference this Kalman-processed dataset.
@@
Figure 10: Counterfactual severing of US–China spillovers. The top panel shows the interaction matrix after zeroing $A_{\text{US},\text{CN}}$ and $A_{\text{CN},\text{US}}$ (highlighted by the black crosses). The bottom panel compares historical GDP (solid), baseline forward simulation (dotted), and the intervention rollout (dash-dot) initialized from the true 2024 state. Removing the bilateral channel slightly dampens post-2024 growth for both economies while leaving other countries largely unchanged, illustrating how targeted coupling edits propagate through the GLV dynamics.

### v2

@@ Results
where $h_{\ell}\in\mathbb{R}^{d_{\ell}}$ denotes the activations at layer $\ell$, $e_{i}\in\mathbb{R}^{d_{e}}$ is the static embedding for country $i$, and $\odot$ is element-wise multiplication. Intuitively, the adapters rescale and shift each hidden unit so that heterogeneous countries can share the same base network while having country-specific gain and bias patterns; this greatly reduces the parameter count relative to training a separate network per economy, yet retains the ability to express qualitatively different responses to shocks. The latent macro network $\phi_{3}$ receives $y_{i},x_{i},e_{i}]$ and outputs $\dot{y}_{i}$, letting unemployment, inflation, and other channels react causally to GDP and other macro channels deviations within the same differentiable rollout. For more details on the mathematical formulation for the forward pass of the macro-economic model, detailing the element-wise dynamics for country $i$ and the FiLM adaptation mechanism, see Appendix Appendix C: Macroeconomic case study.
@@
The interaction term $\psi$ reduces to the classical GLV form $x_{i}(\max(x_{j},0)+\varepsilon)^{\beta}$ when adapters are neutral. We learn $\beta$ jointly with $A$ (converging to 0.20.20.40.4 in practice) to attenuate spillovers from very large economies, and optionally gate each row with $\tanh(\omega_{i})$ when interaction scaling is enabled. The matrix $A$ starts from a weakly coupled baseline ($A_{ij}=1/N$, $A_{ii}=0$).
@@
Training on 1995–2021 data followed by 2021–2024 evaluation shows that the shared-parameter GLV reproduces the GDP trajectories in-sample, while learning interpretable couplings, as illustrated in Fig. 3. Note, that the inferred interaction matrix $A$ for GLV model is not symmetric e.g. China profits more from US GDP than US profits from China GDP, while US and Japan are in mutualism relationship.
We implement counterfactual edits such as zeroing the bilateral US–China channels $A_{\text{US},\text{CN}}$ and $A_{\text{CN},\text{US}}$, and analyze the impact on the global economy. In the appendix Fig. 10, we show the simulation until 2028 with baseline forecasts, and the edited intervention rollout initialized from the true 2024 state, highlighting how severing a single pair of links modifies the GDP trajectories for the US and China, and the global economy. One can observe that all global economies go down, but the largest effect is on the US and China. Note, that the roll-out of 4-year ahead is only used to illustrate the counterfactual effect, see Section Policy implications and Ethics Statement for the policy implications and ethics statement.
@@ Conclusion
Empirically, we showed that a model trained on a single trajectory generalizes out-of-sample to unseen graphs, accurately reproducing aggregate epidemic curves. We further demonstrated intervention analysis: despite never observing interventions during training, the learned model reliably predicts the system’s response, with predicted curves closely matching the ground truth. In particular, we showed that the model accurately captures the response to a time-bounded social-distancing window, highlighting its utility for counterfactual scenarios and control [Bottcher2022AIPontryagin]. The key reason these models succeed in counterfactual analysis is that they learn the underlying dynamics of the system rather than mere correlations, enabling them to predict responses to novel interventions. However, the model cannot respond to arbitrary interventions; it is limited to the interventions that can be incorporated into the learned differential equations, via additive or multiplicative terms in the differential equations, and is not fundamentally changing the underlying dynamics. We have compared the model to other graph neural networks in neural ODE setting, and showed that abm-informed neural networks can better capture the dynamics out of sample than the state-of-the-art graph neural networks. Furthermore, abm-nn model is robust to small perturbations in the data, such as observational noise, and even, it can act as a regularization term for the stability of the model. Finally, we have applied the model to the macroeconomic case study, and showed that it can learn interpretable couplings between macro variables and GDP, and can be used to simulate counterfactual scenarios by editing the interaction structure.
By moving away from the theoretical purity of perfect rationality and equilibrium assumptions, our framework better reflects the reality of complex systems, where agents are not perfectly rational and the system is not operating in equilibrium. We enforce informational boundedness via restricted Graph Neural Networks, ensuring agents react only to local peers rather than global aggregates. Furthermore, the Neural ODE framework allows us to model bounded rationality via several mechanisms: solver step size, optimizer properties, and the complexity of the neural network architecture.
@@
The discretization step of the underlying solver acts as a proxy for cognitive latency, where the step size dictates how quickly agents can process and react to messages from their neighbors. This temporal constraint is matched by a computational one; the use of gradient descent calibrated against empirical data implies that agents learn ’good enough’ local optima—heuristics—rather than global solutions, mirroring human satisficing behavior.
While our framework allows each agent to possess unique decision-making logic (distinct weights)—an approach that works well in data-rich, simulated environments—real-world economic scenarios are often data-scarce. Therefore, we prioritize model parsimony by employing a shared feed-forward neural network modulated by Feature-wise Linear Modulation (FiLM) adapters for each agent.
By combining this FiLM-based heterogeneity with restricted graph topologies, we effectively replace the frictionless ’representative agent’ with a population of diverse, friction-constrained actors. Ultimately, by integrating these temporal, computational limits with heterogeneity, conservation laws, the model offers a plausible mechanism for simulating how constrained agents navigate structural changes.
@@
To conclude, we believe that the proposed framework is not a final solution, but rather a starting point for principled, interpretable neural modeling of complex systems. Future work will explore stochastic differential equation (SDE) formulations to explicitly model process noise and uncertainty quantification, extended Kalman filtering applications, online learning of the dynamical rules, as well as scaling to larger heterogeneous agent populations for more realistic modelling.
Abbreviations
ABM: Agent-Based Model / Agent-Based Modeling
ABM-NN: Agent-Based-Model informed Neural Network
CPI: Consumer Price Index
ETO: Earth-Trisolaris Organization (fictional [Liu2008ThreeBody])
FFN: Feed-Forward Network
FiLM: Feature-wise Linear Modulation
FRED: Federal Reserve Economic Data
GCN: Graph Convolutional Network
GDP: Gross Domestic Product
GLV: Generalized Lotka-Volterra
GNN: Graph Neural Network
GNVF: Graph Neural Vector Field
GraphGPS: Graph General Powerful Scalable Transformer
GraphSAGE: Graph Sample and Aggregate
HNN: Hamiltonian Neural Network
KMC: Kinetic Monte Carlo
MAPE: Mean Absolute Percentage Error
MLP: Multi-Layer Perceptron
NN: Neural Network
ODE: Ordinary Differential Equation
OOS: Out-of-sample
PDE: Partial Differential Equation
QABM: Quantitative Agent-Based Model
ReLU: Rectified Linear Unit
RNN: Recurrent Neural Network
SDE: Stochastic Differential Equation
SIR: Susceptible-Infected-Recovered
@@
The models presented in this paper fall within the category of agent-based models coupled with neural networks. If these models are used to inform policy recommendations, they should be evaluated with extra due diligence, where all assumptions are explicitly reported together with their limitations. Models that are evaluated only in simulation environments have inherent limitations that arise from the fidelity and scope of the simulation environment itself. Real-world deployment requires careful validation against empirical data and consideration of factors not captured in the simulation. Direct interpretation of Generalized Lotka-Volterra macroeconomic model should be done with caution, as it is a simplified model of the real world dynamics, and implications can have serious consequences, especially in the case of global economic tensions. Finally, note, that real-world systems are self-adaptive, and it may not be trivial to separate the effect of the intervention to learned dynamics from the self-adaptive response adaptation of the dynamical rules $\phi_{x},\phi_{y},\psi\rightarrow\phi_{x}^{\prime},\phi_{y}^{\prime},\psi^{\prime}$. For more realistic applications, one should assume that the dynamical rules are adapting in an online manner.
Any model of socio-techno-economic systems, should not be used in unethical ways to harm people, countries, or organizations. Practitioners deploying these models must consider potential misuse, including discriminatory applications, manipulation of social systems, or deployment without informed consent of affected populations. Models of social systems should be designed with the principle in mind that they increase the well-being of humans. Decision-makers must ensure that model-driven interventions respect democratic principles, protect human rights, and account for distributional equity and justice.
@@ Acknowledgements
The experiment in Fig. 4 uses the following hyperparameters: $\text{epochs}=500$, base learning rate $5\times 10^{-4}$ (AdamW with cyclical max learning rate $2\times 10^{-2}$, cyclical step size =75=75), RK4 rollout horizon $t_{\max}=250$ with training truncated at $t_{\text{end}}=50$, curriculum initialized at $T_{\text{base}}=5$ and incremented by $4$ every $50$ epochs, gradient clipping at $2.0$, and stagnation threshold $50$ epochs before forcing a curriculum increase.
@@
The interaction matrix $A$ encodes strategic relationships between factions. The diagonal elements $A_{ii}=-0.001$ represent self-limiting effects (negative self-interaction). Off-diagonal elements capture asymmetric interactions: $A_{\text{Fleet},\text{Earth}}=+0.004$ (the Fleet benefits from Earth’s economy), $A_{\text{Earth},\text{Fleet}}=-0.005$ (Earth is drained by Fleet investment), $A_{\text{Fleet},\text{ETO}}=-0.01$ and $A_{\text{Wallfacers},\text{ETO}}=-0.015$ (both the Fleet and Wallfacers are actively opposed by the ETO). The matrix is sparse, with most entries zero, reflecting that not all factions directly interact in the narrative.
@@
We assess how the macro-level learner behaves when the aggregate SIR curves are corrupted by observational noise. Using the noise-robustness variant of the learning algorithm, we replay the SIR-on-graph experiment with the same Erdős–Rényi prior ($n=100$, $p=0.05$), base parameters ($\beta=0.4$, $\gamma=0.2$), curriculum schedule (1.0 horizon increment every 50 epochs starting at $t=1$), and the triangular learning-rate policy ($\text{base }lr=10^{-4}$, $\text{max }lr=5\times 10^{-4}$, step 100). We vary only the observational noise injected into $s(t),i(t),r(t)$ for $t\geq 3$: Gaussian noise with $\sigma\in\{1.0,2.0,3.0\}$ is added per time step, while the differentiable learner sees only these noisy trajectories. Each run trains for 1000–1500 epochs (longer for higher noise).
Figure 6: Noise-robustness study for the macro SIR learner. Columns correspond to observational noise $\sigma=1.0$ (up), $2.0$ (middle), and $3.0$ (down). Each panel overlays the true aggregate curves (solid), noisy observations (dotted), and learned trajectories (dashed with diamonds). Despite the increasing corruption after $t=3$, the ABM-NN continues to track the underlying dynamics and maintains separation between in-sample (denoted by the shaded training region) and out-of-sample forecasts.
@@ 3. Detailed Component Definitions
$\psi_{i}=\mathcal{C}(k)\cdot[x_{i}]_{+}\sum_{j}A_{ij}([x_{j}]_{+}+\epsilon)^{\beta},$ (44)
@@
The optimizer learns the interaction matrix $\mathbf{A}$, the exponent $\beta$.
@@
    • FiLM adapters applied to all hidden layers.
@@
    • Training period: 1995–2022 (cutoff year).
@@
The macro panel used throughout the paper originates from a data utility pipeline, which queries the World Bank API for the ten largest economies (GDP, unemployment, lending interest rate, sovereign debt, working-age population share, CPI inflation, current account balance) and augments the table with the global oil price index from FRED (POILAPSPINDEXQ). We then apply a strictly causal constant-velocity Kalman filter to each country/indicator pair using variable-specific process/measurement noise levels; missing values are imputed forward in time without peeking at future observations and the result is stored in the processed dataset.
@@
Figure 10: Counterfactual severing of US–China spillovers. The top panel shows the interaction matrix after zeroing $A_{\text{US},\text{CN}}$ and $A_{\text{CN},\text{US}}$ (highlighted by the black crosses). The bottom panel compares historical GDP (solid), baseline forward simulation (dotted), and the intervention rollout (dash-dot) initialized from the true 2024 state. Removing the bilateral channel (between US and China) dampens post-2024 growth for both economies and lowers the global economy growth for the rest of the world, illustrating how targeted coupling edits propagate through the macroeconomic dynamics.